OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 14 vowels and 22 consonants, what is the probability you will choose a vowel then a consonant ?

OpenStudy (anonymous):

@hartnn @haitianmontana

OpenStudy (anonymous):

@Salmon are you able to teach me how this is done?

OpenStudy (anonymous):

In total their are 36 letters right??

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

First we will find our sample space In how many ways you can select the first letter??

OpenStudy (anonymous):

14/36 because they want a vowel first.

OpenStudy (anonymous):

Ok...you are actually approaching the answer in different way but even your method is right.. Lets proceed in your way... Now find the probability of choosing a consonant...

OpenStudy (anonymous):

22/36

OpenStudy (anonymous):

now the probability that you will choose a vowel then a consonant is the product of their individual probabilities i.e probability that you will choose a vowel then a consonant=(probability of chossing a vowel)(probability of choosing a consonant)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What is the probability of chossing a vowel?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

you mean 14/36?

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

probability of choosing a consonant= 22/36 so your answer is (14/36)(22/36)=???

OpenStudy (anonymous):

idk 308/36????

OpenStudy (anonymous):

\[\frac{ 14 }{ 36 }\times \frac{ 22 }{ 36 }=\frac{ 14\times22 }{ 36\times36 }=\frac{ 308 }{ 1296 }\]

OpenStudy (anonymous):

i tried that it doesn't work....

OpenStudy (anonymous):

either that or i can't simplify... but whatever i give up

OpenStudy (anonymous):

k..lets start again....This time in an systematic way...